OpenStudy (anonymous):

We are studying elastic collisions. There is this system of equations problem where we are solving for the first and second final velocities. Help! equations: m1v1,i+m2v2i = m1v1f+m2v2f 0.5m1v1i^2+0.5m2v2i^2 = 0.5m1v1f^2 + 0.5m1v1f^2 Givens: m1= 2kg m2=3kg v1,i= 4m/s v2,i=-5m/s unknown: v1,f= ? v2,f=? Please help!

OpenStudy (shamim):

So u hv 2 equations and u hv 2 unknown parameter.

OpenStudy (shamim):

So its possible to find out v1f and v2f

OpenStudy (shamim):

After putting ur given values we can write 2*4+3*5=2*v1f+3*v2f 8+15=2*v1f+3*v2f 23=2*v1f+3*v2f..........(1)

OpenStudy (shamim):

Now ur second equation 0.5*2*4^2+0.5*3*5^2=0.5*2*v1f^2+0.5*3*v2f^2 16+37.5=v1f^2+1.5*v2f^2.........(2)

OpenStudy (shamim):

Now these 2 equtions for getting the value of v1f=? v2f=?

OpenStudy (shamim):

Response plz!!!!!!

OpenStudy (anonymous):

Sorry, I was gone basically it means m1v1i+m2v2i = m1vlf+m2v2i we dont know the final and initial velocities but if you plug it in for both we can find the other two. But I get it now it was simple enough. Since we have only two things we dont know and three things we know it is solvable.